HUBBLE SPACE TELESCOPE DAILY REPORT #1061 PERIOD COVERED: 0700 EST Daily Status Report as of 026/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 4721 (WF/PC -2 Decontamination and TEC Turn On) The WF/PC-2 was used to make ultraviolet photometric observations of the star standard A+81D266, plus K-Spots, and dark calibration images. All four sets of CCDs were used. A total of 22 images (i.e., two sets of 11 each) were received and displayed in the OSS. No anomalies were reported. 1.2 Completed Two Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Fields Calibration) The WF/PC-2 was used to make a set of visual internal flat images. 17 images were received and displayed. Four of the images were reported to be underexposed, and the other images were okay. 1.3 Completed Two Sets GHRS/2 4731 (COSTAR/GHRS Coarse Alignment) The High Resolution Spectrograph was used to make observations of the ultraviolet standard star AGK81D266 to help refine the tilt and focus of the COSTAR mirrors. The mirror motions were successfully uplinked to the spacecraft. The onboard target acquisition was successful, and left the target star well centered. The subsequent observations were made in fine lock. No anomalies were noted. 1.4 Completed WF/PC-2 4819 (WF/PC-2 PSF Characterization) The WF/PC-2 was used to make multiple exposures of the globular cluster Omega Centauri during coarse and fine dwell scans. The fine scans were only 0.025 arcseconds apart. Two different filters were used. 45 observations were received and displayed. The images were reported to look normal. 1.5 Completed FOC/96 5151 (Follow-up Far-UV Objective Prism Observations of High Red Shifted Objects) The Faint Object Camera (f/96) was scheduled to make ultraviolet objective prism images of the quasi-stellar object Q0302-003 to determine if the object exhibits intergalactic He II absorption lines. The guide star acquisition failed to reach fine lock, and FGS#2 had no guide star at all. FGS#1 failed to coarse track. No observations were made. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 5 Scheduled Reacquisitions : 4 Successful : 4 At 026/0908Z the FGS acquisition failed during Loss Of Signal (LOS). FGS#1 failed to course track. (See Observational Report 1.5) 2.2 FHST Updates: Scheduled :14 Successful :14 2.3 Significant HSTARs: None 2.4 Operations Notes: All HST subsystems are operating normally, as SMOV activities continue to be successfully performed in accordance with the planned timeline. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium High Gal. Lat) 3.2 Complete WF/PC-2 5240 (SMOV-WFPC-2 Flat Fields Calibration) 3.3 Complete Two Sets of 4731 (COSTAR/GHRS Coarse Alignment) 3.4 Complete WF/PC-2 5475 (A Snapshot Survey for Blue Stragglers in Magellanic Cluster Cores) 3.5 Complete WF/PC-2 5590 (Cycle 4 DD HST Observations of Comet Shoemaker-Levy (1993E)) 3.6 Complete WF/PC-2 4749 (WF/PC-2 Relative Photometric Calibration) 3.7 Complete FGS 5583 (Long Term Stability of FGS#3) 3.8 Complete FOS/RD/COSTAR 5605 (COSTAR/FOS Fine Alignment - X-Tilt/Y- Tilt Repeat of 4747) 3.9 Complete S/C 5254 (GHRS SMOV Side 1 Activation) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: None /DMH -END-